13. + . Solve: 2. Suppose: 21 = 2+1, 2=3–21, and 23=- (a) 321 +421 (b) +-3 +421-8
Answer #1

Z, = 2+i Z,= 3 -२ 32+५३, = 3ai) + ५ (3-21) 6 +31 t12-8i = ।४ -5। 132,+4831 (६}+5 %3D 3५१ %3D 3 -32 + ५६-४ ৬ - 1}} = 43 4 323